### v1

# Beyond Full Label: Single-Point Prompt for Infrared Small Target Label Generation

Abstract
In this work, we make the first attempt to construct a learning-based single-point annotation paradigm for infrared small target label generation (IRSTLG). Our intuition is that label generation requires just one more point prompt than target detection: IRSTLG can be regarded as an infrared small target detection (IRSTD) task with the target location hint. Based on this insight, we introduce an energy double guided single-point prompt (EDGSP) [^1] Code will be made public at https://github.com/xdFai. framework, which adeptly transforms the target detection network into a refined label generation method. Specifically, the proposed EDGSP includes: 1) target energy initialization (TEI) to create a foundational outline for sufficient shape evolution of pseudo label, 2) double prompt embedding (DPE) for rapid localization of interested regions and reinforcement of individual differences to avoid label adhesion, and 3) bounding box-based matching (BBM) to eliminate false alarms. Experimental results show that pseudo labels generated by three baselines equipped with EDGSP achieve 100% object-level probability of detection (${{P}_{d}}$) and 0% false-alarm rate (${{F}_{a}}$) on SIRST, NUDT-SIRST, and IRSTD-1k datasets, with a pixel-level intersection over union (IoU) improvement of 13.28% over state-of-the-art label generation methods. Additionally, the downstream detection task reveals that our centroid-annotated pseudo labels surpass full labels, even with coarse single-point annotations, it still achieves 99.5% performance of full labeling.
Introduction
Single-frame infrared small target detection (IRSTD) plays a crucial role in maritime rescue, early warning, and high-precision navigation (Wang, Zhou, and Wang 2019; Kou et al. 2023a, b; Zhang et al. 2024b). To isolate small targets from complex clutter, high-performance IRSTD methods typically extract discriminative features from extensive paired infrared images and pixel-level masks (Zhang et al. 2022; Wu et al. 2023; Chen et al. 2024a). This learning paradigm highlights the need for large-scale, accurate mask annotation. However, infrared (IR) small targets typically occupy only a few pixels, exhibit a low signal-to-clutter ratio, and lack well-defined edges. These characteristics render manual pixel-by-pixel annotation time-consuming and error-prone. Consequently, an efficient and reliable infrared small target label generation (IRSTLG) method is urgently required.
Figure 1: Task correlation between (a) single-point supervised IRSTD, (b) fully supervised IRSTD, (c) model-based IRSTLG, and (d) proposed learning-based IRSTLG.
Single-point annotation [^2] Unless explicitly stated, centroid single-point is used. (SA) has recently demonstrated efficiency in labeling IR small targets, reducing the time needed by 80% compared to pixel-level annotation. As shown in Figure 1(a), LESPS (Ying et al. 2023) made the first attempt to adopt SA for weakly supervised IRSTD. The mapping degradation is cleverly introduced to extract the target contour during network training. Then, these intermediate predictions are employed to guide the point label evolution. Precisely, LESPS integrates IRSTD network training with label generation into a unified framework. Subsequently, a standardized label generation paradigm emerged. As illustrated in Figure 1(c), these pioneering works use SA to guide traditional model optimization for generating pseudo masks, such as MCLC (Li et al. 2023a), MCGC (Kou et al. 2024), and COM (Li et al. 2024). Note that while these three works are claimed as “Single-Point Supervision”, they fundamentally belong to the label generation task with the single-point prompt as input. To avoid confusion, Figure 1(a)-(c) elucidates two critical links between these tasks. 1) If replace the full label in IRSTD with the pseudo label obtained in IRSTLG, single-point supervised IRSTD can be treated as an organic combination of single-point prompt IRSTLG and fully supervised IRSTD. This implies that developing IRSTLG can directly improve point-supervised IRSTD. 2) Label generation requires just one more point prompt than detection. Thus, IRSTLG can be regarded as an infrared small target detection task with target location hint. This indicates that the labeling task adheres to at least two essential criteria.
  • The pseudo label should satisfy accurate annotation (AA), i.e, object-level ${{P}_{d}}$=100%, ${{F}_{a}}$=0%.
  • IRSTLG frees up more effort to specialize in target contour than IRSTD, therefore should have higher IoU.
To validate this reasoning, Figure 2 showcases a comparison between the detection results of five IRSTD methods (Wu et al. 2023; Hou et al. 2022; Li et al. 2022; Wu, Hong, and Chanussot 2022; Yuan et al. 2024) with the pseudo masks generated by three IRSTLG algorithms (Li et al. 2023a; Kou et al. 2024; Li et al. 2024). Experimental results indicate that existing IRSTLG methods fail to meet standards in AA and IoU. This issue arises because these traditional algorithms neglect to leverage target location cues to extract features from fully labeled data.
Since well-designed IRSTD networks already achieve satisfactory target shape inference, and single-point prompts can be easily extracted from full labels, the primary goal of this paper is to design an effective “Upgrade Tool” that transforms the IRSTD network into a label generation method using the single-point prompt. As shown in Figure 1(d), the essence of this paradigm is learning point prompt-based mask labeling experience from already labeled data. In fact, point prompt combined with segmentation baseline has been applied to interactive annotation for general objects (Wang et al. 2019; Zhai et al. 2023). A common approach involves using multiple extreme points or corners to precisely crop the target region, then pairing it with the encoded prompt and full labels for fully supervised training. Although this pipeline has yielded satisfactory results in routine areas, directly applying it to our task reveals three pitfalls. 1) Missing edge coordinates. Only one pixel of the small target is annotated, the lack of edge coordinates hinders deep models focusing on the target contour. Even with multiple cropping scales and carefully crafted prompts, it still encounters obstacles in achieving accurate annotation (detailed see Experiments). 2) Unique target characteristics. Infrared small targets occupy only 0.15% of the total image pixels, and their edges are blurred by diffraction effects (Kou et al. 2023a). Two small targets in close proximity are easy to stick together and be mistakenly labeled as a single target, reducing ${{P}_{d}}$. 3) Strong background clutter. Cloud edges, sea phosphorescence, and bright buildings in IR images have similar energy distributions to small targets (Zhang et al. 2024a). Even under effective location cues, this similarity still causes the deep model to misclassify pixels in these interference regions, resulting in false alarms.
Figure 2: Comparing IRSTD and IRSTLG methods on mixed SIRST, NUDT-SIRST, and IRSTD-1k in $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{a}}({10}^{-6})$. EDGSP-Coa. denotes EDGSP with the coarse single-point prompt.
To tackle these problems, we propose a novel framework for IRSTLG, named energy double guided single-point prompt (EDGSP). Without requiring complex cropping, EDGSP enables end-to-end model target contours through the integration of local and global context provided by the point prompt and deep model. As shown in Figure 3, this framework consists of three key components. 1) Target energy initialization (TEI): Inspired by IR targets that always have similar energy and shape distribution, we expand each point prompt into a Gaussian with local salience priors as the initial pattern of every target to evolve more complete outlines. 2) Double prompt embedding (DPE): The encoded prompt is embedded twice. The first embedding ensures the model directly focuses on the interested region, even in shallow layers. The second embedding refines the target salience map, enhancing target boundaries and preventing shape merging. 3) Bounding box-based matching (BBM): As a post-processing approach, we generate bounding boxes for each candidate target, then match the point prompt with each box by coordinates to eliminate clutter.
Benefiting from the above structures, as shown in Figure 2, our annotation method showcases powerful shape prediction capabilities and achieves ${{P}_{d}}$=100%, ${{F}_{a}}$=0on three datasets. The main contributions are summarised as follows.
  • To the best of our knowledge, we present the first study of the learning-based IRSTLG paradigm, and introduce EDGSP creating a crucial link between label generation and target detection task.
  • We propose the target energy initialization (TEI), double prompt embedding (DPE), and bounding box-based matching (BBM) strategies to address the insufficient label shape evolution, label adhesion, and false alarm.
  • For the first time, three baselines equipped with EDGSP achieve accurate annotation on three datasets. The downstream detection task illustrates that our pseudo label surpasses the full label. Even with coarse point annotated, EDGSP achieves 99.5% performance of full labeling.
Related Work
Infrared Small Target Detection
IRSTD techniques have been developed over decades, and for a long period, they relied on handcrafted traditional methods such as image filtering (Bai and Zhou 2010; Deshpande et al. 1999), human visual system (Kim and Lee 2012; Chen et al. 2013), and low-rank optimization (Gao et al. 2013; Liu et al. 2023). Recently, deep learning-based methods have been able to obtain high-level semantics from large amounts of paired data by stacking nonlinear feature extractors, achieving robust detection in complex scenes. Since pixel-by-pixel classification is crucial for subsequent target recognition, segmentation-based IRSTD is a key emphasis in these data-driven approaches. Based on a U-shaped structure (Ronneberger, Fischer, and Brox 2015), ACM (Dai et al. 2021) improves the feature fusion strategy of high-level semantics with low-level details using an asymmetric context modulator. To mitigate the semantic gap between encoders and decoders, DNA-Net (Li et al. 2022) adopts dense nested interactive modules to facilitate progressive interaction of cross-scale features and adaptive enhancement of fine-grained features, while UIU-Net (Wu, Hong, and Chanussot 2022) directly nest multiple small U-Net into each codec of U-shaped structure. To address the limitations in long-range information modeling, MTU-Net (Wu et al. 2023) employs a self-attention to compute the spatial correlations of fused features, thereby constructing a global representation of the image. Most recently, SCTransNet (Yuan et al. 2024) explicitly interacts and enhances full-level encoder features by spatial-channel cross transformer blocks to effectively suppress interference.
In summary, existing research emphasizes designing tight structures to achieve high performance. However, the labor-intensive task of pixel-by-pixel annotation for infrared small targets hampers the advancement of these data-dependent methods. Therefore, this work aims to alleviate the burden of infrared small target labeling.
Point Supervised Segmentation
One working mode related to us is point-supervised segmentation. As a weak supervision form, point supervision only requires one or a few points as label (Qian et al. 2019; Zhao and Yin 2020; Li et al. 2023b). Bearman et al. (Bearman et al. 2016) propose the first single-point supervised semantic segmentation method by integrating an objectiveness prior into a loss function, to elevate segmentation effectiveness. Subsequently, Zhang et al. (Zhang et al. 2023) leverage the common characteristics of target regions and present a contrast-based variational model for point-supervised histopathology image segmentation, achieving more regionally consistent and smoother boundaries. Close to our work, LESPS (Ying et al. 2023) is the first to introduce single-point supervision to infrared small target detection task, based on mapping degradation phenomenon to guide point label evolution.
Unlike the aforementioned works, we still rely on fully annotated masks for supervision, yet employ single-point as a form of input to guide the IRSTD network. This approach reveals how to utilize additional prompts to achieve high-precision pseudo-label generation for small targets.
Infrared Small Target Label Generation
To the best of our knowledge, there are only three IRSTLG efforts. They are all based on the model-based single-point prompt methods. As a pioneer, MCLC (Li et al. 2023a) makes the first attempt to employ a Monte Carlo linear clustering process with random noise and averages the clustering results to obtain reliable pseudo masks. Subsequently, MCGC adaptive recovers small targets with multi-scale chain growth clustering, further enhancing annotation robustness (Kou et al. 2024). Recently, Li et al. (Li et al. 2024) designed a variational level set with an expected difference energy function to address the excessive regularization causing target contours to disappear. However, these works tend to treat the points as a ”Supervision” signal. This biased task orientation causes them to overlook the potential of combining point prompt-derived location cues to extract features from complete labeled datasets, culminating in pseudo labels with unclear edges, target adhesion, and false alarms.
In this paper, we accurately define the operating paradigm where images and single-point prompts serve as inputs and produce pseudo labels as “single-point prompt for infrared small target label generation”. Unlike previous methods, we present the first study of a learning-based paradigm for IRSTLG, which artfully evolves the target detection network into an advanced label generation method with the aid of the single-point prompt.
Figure 3: Illustration of the proposed energy double guided single-point prompt (EDGSP) framework. It consists of three parts. (a) Target energy initialization (TEI) for sufficient shape evolution, (b) double prompt embedding (DPE) to reinforce individual differences and prevent mask adhesion, and (c) bounding box-based matching (BBM) to eliminate false alarms.
Point Prompt Interactive Segmentation
The technique most similar to ours is point prompt interactive segmentation (Rother, Kolmogorov, and Blake 2004; Xu et al. 2016; Liew et al. 2017), which aims to reduce annotation time. DEXTR (Maninis et al. 2018) uses four extreme points for segmentation, the interest region is cropped and the Gaussian heat map is utilized as the encoding form of prompts. Similarly, IOG (Zhang et al. 2020) uses the approximate center of the target as the inside point and a set of symmetrical corner points as the outside points. After cropping the target of interest, the image, along with heat maps is fed into the segmentation model. Luo et al. (Luo et al. 2021) uses the exponentialized geodesic distance transform to convert medical images into Euclidean distance maps for improved foreground-background separation. Recently, SAM (Kirillov et al. 2023) has gained recognition for its flexibility and generalization capabilities. It is a transformer-based model for semantic segmentation that utilizes various types of prompts to encode key areas.
Despite the success of the above work on generic target segmentation, the unique characteristics and annotation patterns of infrared small targets render these strategies ineffective. The proposed EDGSP differs from them in three ways: 1) It processes the entire image directly without cropping. 2) It assigns explicit physical meaning to Gaussian heat maps, thereby using the double embedding strategy. 3) EDGSP introduces a box-based matching approach specifically designed to eliminate false alarms.
Proposed Method
The overview of our proposed EDGSP is shown in Figure 3. It is very concise and consists of three parts.
Target Energy Initialization
The long-range imaging mechanism of infrared systems brings severe diffraction effects, resulting in small targets appearing with a nearly Gaussian distribution in infrared images (Dai et al. 2023). By combining this physical phenomenon with the local contrast prior of small targets (Chen et al. 2013), we propose target energy initialization to predefine a basic profile for each target. As shown in Figure 3, given an input image ${\mathbf{{I}_{i}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$ and corresponding single-point prompt ${\mathbf{{I}_{p}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$ (Since the single-point prompt is a pixel within the target, it can be derived from the already labeled ground truth ${\mathbf{{G}_{t}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$). 2D Gaussian function is used to expand each point in ${\mathbf{{I}_{p}}}$ as an energy blob, thus creating a single init shape map ${\mathbf{{I}_{g}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$. This expansion simplifies the task by shifting from shaping area targets with points to refining contours using Gaussian blobs, aligning the prior distribution of the targets, and aiding the model in learning complete shape representations.
Double Prompt Embedding
Without loss of generality, given a U-shaped IRSTD network $\mathcal{G}$, encoded prompt ${\mathbf{{I}_{g}}}$ embed into $\mathcal{G}$ twice, as shown in Figure 3(b). Considering that edge information can guide shape construction (Zhang et al. 2022), the Sobel operator $\mathcal{S}$ is first applied to the original infrared image ${\mathbf{{I}_{i}}}$ to obtain a rough edge map ${\mathbf{{I}_{s}}}$. Then, ${\mathbf{{I}_{i}}}$, ${\mathbf{{I}_{s}}}$, and the init shape map ${\mathbf{{I}_{g}}}$ are concatenated among the channel dimension as a three-channel input image ${\mathbf{{I}_{t}}}\in\mathbb{R}^{{3}\times{H}\times{W}}$, and get the penultimate layer features ${\mathbf{L}}$ of $\mathcal{G}$ as follows:
${\mathbf{{I}_{t}}}=Concat({\mathbf{{I}_{i}}},\mathcal{S}({\mathbf{{I}_{i}}}),{\mathbf{{I}_{g}}})$ (1)
${\mathbf{L}}={\mathcal{G}^{Conv}_{w/o}}({\mathbf{{I}_{t}}})$ (2)
where ${\mathcal{G}^{Conv}_{w/o}}$ stands for $\mathcal{G}$ that without the last convolution layer. In the second embedding, ${\mathbf{{I}_{g}}}$ is treated as a semantic map that reinforces the target quantity, it is concatenated with the ${\mathbf{L}}$ along the channel dimension to refine the model’s final mapping and prevent close-range target sticking.
${\mathbf{{O}_{s}}}={Sigmoid}({Conv}(Concat({\mathbf{L}},{\mathbf{{I}_{g}}})))$ (3)
where ${\mathbf{{O}_{s}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$ represents the saliency map, ${Conv}$ denotes the last layer of $\mathcal{G}$, typically constructed with a $1\times 1$ or $3\times 3$ convolution that squeezes channel and preserves the spatial dimensions. The binary cross entropy loss is then used to compute pixel-wise differences between the ${\mathbf{{O}_{s}}}$ and the ground truth ${\mathbf{{G}_{t}}}$.
Table 1: Comparison with SOTA IRSTLG methods on SIRST, NUDT-SIRST, IRSTD-1k in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{at}}(\%)$, ${{F}_{a}}({10}^{-6})$. The best and the second-best results are highlighted in bold and underlined, respectively. EDGSP-Coa. indicates the EDGSP under the coarse single-point prompt.
Method SIRST NUDT-SIRST IRSTD-1k Total
$IoU\uparrow$ ${{P}_{d}}\uparrow$ ${{F}_{at}}\downarrow$ ${{F}_{a}}\downarrow$ $IoU\uparrow$ ${{P}_{d}}\uparrow$ ${{F}_{at}}\downarrow$ ${{F}_{a}}\downarrow$ $IoU\uparrow$ ${{P}_{d}}\uparrow$ ${{F}_{at}}\downarrow$ ${{F}_{a}}\downarrow$ $IoU\uparrow$ ${{P}_{d}}\uparrow$ ${{F}_{at}}\downarrow$ ${{F}_{a}}\downarrow$
LESPS 41.84 97.71 1.52 0.27 38.86 98.94 0.95 2.06 38.76 98.98 0.68 0.04 39.28 98.73 1.00 0.14
COM 53.19 83.65 10.3 1.85 8.026 52.54 32.1 6.96 41.21 69.39 20.1 1.12 13.84 61.29 25.9 3.51
MCLC 75.63 98.48 3.80 0.68 65.08 95.87 12.2 2.64 69.60 98.64 4.42 0.25 67.98 96.87 9.19 1.25
MCGC 72.43 97.33 0.76 0.14 77.67 99.68 0.32 0.07 63.24 99.66 0.34 0.02 72.80 99.26 0.40 0.05
SAM 6.273 100.0 0.76 0.14 8.466 100.0 0.74 0.16 2.428 100.0 1.02 0.06 6.41 100.0 0.80 0.11
Robust SAM 78.89 98.86 0.38 0.07 67.05 99.79 0.21 0.05 72.28 99.32 0.34 0.02 70.14 99.53 0.27 0.04
EDGSP-Coa. 83.05 99.62 0 0 94.23 99.89 0.11 1.01 71.07 99.66 0.34 0.47 85.36 99.80 0.13 0.62
EDGSP 83.83 100.0 0 0 95.51 100.0 0 0 73.80 100.0 0 0 87.07 100.0 0 0
Bounding Box-based Matching
In previous works, eight connective regions matching is popular to eliminate false alarms during the label evolution process (Ying et al. 2023; Kou et al. 2024), ensuring that evolution information consistently applies to the initial point. However, this strategy ignores cases where the point prompt is located on or outside the target boundary, leading to incorrect processing results [^3] The shortcomings of this approach are detailed in the supplementary material.. We tackle this issue by adding a compact design. In Figure 3(c), after performing eight-neighborhood clustering on the threshold-segmented candidate map ${\mathbf{{O}_{c}}}$, we generate bounding boxes for each component to create a b-box map ${\mathbf{{B}_{m}}}$. Then, match the point prompt ${\mathbf{{I}_{p}}}$ with ${\mathbf{{B}_{m}}}$. Specifically, if the point in ${\mathbf{{I}_{p}}}$ falls inside a bounding box, the target in this box is retained; otherwise, it is discarded as a false positive. After performing this region match, we got the final pseudo label ${\mathbf{{O}_{p}}}$.
Experiments
Experimental Setup
Datasets. Experiments are conducted on three mixed datasets: SIRST (Dai et al. 2021), NUDT-SIRST (Li et al. 2022), and IRSTD-1k (Zhang et al. 2022), which include 427, 1327, and 1001 infrared images with one or more targets, respectively. We used the method from (Li et al. 2022) to partition the datasets for SIRST and NUDT-SIRST, and followed (Zhang et al. 2022) for splitting IRSTD-1k.
Figure 4: Visual results of seven IRSTLG methods on the SIRST, NUDT-SIRST, IRSTD-1k. EDGSP-Coa. denotes the pseudo label of EDGSP under coarse single-point prompt. Small flaws in ground truth are highlighted with arrows. Refer to supplementary material for more visual results.
Implement Details. SCTransNet (Yuan et al. 2024) is used as the backbone of EDGSP. The peak half-width of the Gaussian is set to 4 and the segmentation threshold is set to 0.5. We compare all IRSTLG methods: MCLC (Li et al. 2023a), MCGC (Kou et al. 2024) and COM (Li et al. 2024), and learning-based single-point annotation methods: SAM (Kirillov et al. 2023) and Robust SAM (Chen et al. 2024b). Furthermore, we also include the point-supervised IRSTD framework LESPS (Ying et al. 2023), using DNA-Net as the backbone, consistent with the original paper. We provide more details of experiments in the supplementary material.
Table 2: The improvement brought by the EDGSP framework to seven SOTA IRSTD methods on SIRST, NUDT-SIRST, IRSTD-1k in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{at}}(\%)$, ${{F}_{a}}({10}^{-6})$, $Flops(G)$, and $Params(M)$.
Method IRSTD Baseline IRSTD with EDGSP
$IoU\uparrow$ ${{P}_{d}}\uparrow$ ${{F}_{at}}\downarrow$ ${{F}_{a}}\downarrow$ ${Flops}\downarrow$ ${Params}\downarrow$ $IoU\uparrow$ ${{P}_{d}}\uparrow$ ${{F}_{at}}\downarrow$ ${{F}_{a}}\downarrow$ ${Flops}\downarrow$ ${Params}\downarrow$
ACM 61.43 94.80 13.75 64.50 0.402 0.398 73.60 99.60 0.066 0.28 0.405 0.398
ALCNet 63.26 94.98 13.69 58.74 0.378 0.427 74.06 99.67 0.066 1.71 0.381 0.427
MTU-Net 71.62 94.15 12.16 28.31 6.194 8.221 84.20 100.0 0 0 6.215 8.221
ISTDU 80.12 96.74 9.103 32.18 7.944 2.752 84.07 99.93 0.066 1.11 7.973 2.752
DNA-Net 80.53 96.90 5.249 15.28 14.26 4.697 86.78 100.0 0 0 14.28 4.697
UIU-Net 82.40 97.34 6.389 15.39 54.43 50.54 86.96 99.93 0 0 54.50 50.54
SCTransNet 83.43 97.28 5.781 8.622 10.12 11.19 87.07 100.0 0 0 10.16 11.19
Average 73.81 95.79 9.993 34.32 13.39 11.18 82.19 99.85 0.041 0.43 13.42 11.18
Difference - - - - - - +8.38 +4.06 ×0.35% ×1.25% +0.04 +0
Evaluation Metrics. To evaluate pseudo labels and downstream target detection results, intersection over union (IoU) is used as a pixel-level metric to assess the model’s contour inference capability (Dai et al. 2021). probability of detection (${{P}_{d}}$) and false-alarm rate (${{F}_{a}}$) are utilized as target-level evaluation metrics to assess localization capabilities. The details of ${{P}_{d}}$, ${{F}_{a}}$ are same as (Li et al. 2022). Considering the number of false alarms is more important for the labeling task, the target false-alarm rate (${{F}_{at}}$) is defined as the ratio of false predicted target numbers ${T}_{false}$ and all targets ${{N}_{all}}$.
Quantitative Results
As shown in Table 1, only EDGSP achieves ${{P}_{d}}$=100%, ${{F}_{a}}$=0%, ${{F}_{at}}$=0across three datasets at the same time. This indicates that our framework is currently the only one ensuring the most basic level of annotation quality. Additionally, EDGSP’s IoU surpasses the state-of-the-art (SOTA) labeling method (MCGC) by 14.27, demonstrating a superior shape prediction capability. Even with the coarse single-point prompt, EDGSP outperforms other methods with the centroid single-point prompt.
Qualitative Results
Figure 4 illustrates that the results of EDGSP are closer to the target silhouette in the infrared image than those of other models. In fact, many pseudo masks such as Misc_28, Misc_74, 000595, XDU104, and XDU697 even exceed the quality of ground truth (GT). We also observe that COM and MCLC tend to merge the target with the background, while LESPS, MCGC, and Robust SAM struggle to accurately infer the morphology of infrared small targets. Additionally, SAM is incapable of labeling infrared small targets.
Figure 5: Quantitative results of three baselines at six cropping scales in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{at}}(\%)$. Dashed and solid lines represent label generation results from direct center-cropping and from using EDGSP on the cropped image, respectively.
Upgrade Tool: EDGSP
As shown in Table 2, with the help of EGDSP, the average IoU and ${{P}_{d}}$ values for seven SOTA IRSTD methods, including ACM (Dai et al. 2021), ALCNet (Dai et al. 2021), MTU-Net (Wu et al. 2023), ISTDU (Hou et al. 2022), DNA-Net (Li et al. 2022), UIU-Net (Wu, Hong, and Chanussot 2022) and SCTransNet (Yuan et al. 2024), improved by 8.38% and 4.06%, respectively, while the ${{F}_{at}}$ decreased by a factor of 243.73. Flops and parameters remained nearly constant. Three IRSTD methods (MTU-Net, DNA-Net, and SCTransNet) achieve accurate annotations. This indicates that EDGSP builds a reliable bridge between IRSTD and IRSTLG. We also found a positive correlation between pseudo-label performance and detection results, indicating that future iterations of the IRSTD method can further enhance pseudo-label quality through EDGSP. Given that SCTransNet achieved the highest IoU score, we selected it as the backbone for EDGSP unless stated differently.
Table 3: Based on SCTransNet (SCT), ablation study of the single prompt embedding (SPE), TEI, rough edge map (REM), DPE, and BBM on SIRST, NUDT-SIRST, and IRSTD-1k. SPE means directly concatenating the single-point prompt with the original image as input. More details of each module are discussed in the supplementary material.
Method SCT SPE TEI REM DPE BBM $IoU\uparrow$ ${{P}_{d}}\uparrow$ ${{F}_{at}}\downarrow$
M1 $\surd$ 83.43 97.28 5.78
M2 $\surd$ $\surd$ 86.16 99.67 1.73
M3 $\surd$ $\surd$ $\surd$ 86.52 100.0 1.74
M4 $\surd$ $\surd$ $\surd$ $\surd$ 86.74 99.93 1.93
M5 $\surd$ $\surd$ $\surd$ $\surd$ 87.02 100.0 1.59
Ours $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ 87.07 100.0 0
Prompt-based Center-Cropping
Figure 5 showcases the results of three baselines (U-Net, DNA-Net, and SCTransNet) under prompt-based center-cropping. With the crop size increasing, model performance initially improves and then declines, and fails to achieve high-performance labeling (${P}_{d}<100\%$, IoU$<$85). This is because small crop sizes fail to encompass large targets, while large crop sizes introduce excessive irrelevant padding at the image edges. Additionally, target overlap is challenging to manage when splicing the inferred image back into the original image. In summary, missing boundary coordinate information diminishes the effectiveness of the formal center-cropping strategy. Therefore, EDGSP directly processes the entire image.
Table 4: Quantitative results of five IRSTD trained with five types of pseudo-labels and full labels on downstream test sets in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{a}}(\%)$. The best and the second-best results are highlighted in bold and underlined, respectively.
Method Pseudo Label Full Label
MCLC MCGC Robust SAM EDGSP-Coa. EDGSP
ACM 61.17/90.07/22.16 61.42/91.72/23.57 62.43/92.38/29.22 65.03/92.38/23.97 65.59/91.39/23.92 65.65/93.38/32.70
RDIAN 64.94/94.04/28.54 68.38/93.38/16.96 65.57/93.38/24.43 73.27/93.05/21.26 74.68/93.71/18.23 72.64/92.72/13.93
MTU-Net 66.33/89.07/18.00 68.24/93.71/10.49 66.54/93.38/20.36 75.46/93.38/9.95 75.31/95.36/17.96 75.89/94.70/13.48
DNA-Net 66.12/93.05/15.29 67.80/93.38/6.60 67.91/94.37/12.48 81.52/94.70/7.55 81.53/96.69/8.14 81.63/95.70/9.00
UIU-Net 66.07/93.05/44.06 69.55/93.71/4.26 66.57/93.70/6.02 81.14/94.70/8.73 82.70/96.03/6.02 82.46/96.68/7.46
Average 64.93/91.86/25.61 67.08/93.18/12.38 65.80/93.44/18.50 75.28/93.64/14.29 75.96/94.64/14.85 75.65/94.64/15.31
Ablation Study
As shown in Table 3, SCTransNet is used as the baseline to validate the effectiveness of each proposed strategy. In M2, just incorporating single prompt embedding, the value of ${P}_{d}$ and IoU exceeds the full cropping strategies in Figure 5, illustrating the necessity of entire image processing. Using SPE, the IoU value further improved, and ${{P}_{d}}$ reached 100%, which demonstrated the SPE is capable of fleshing out target shapes and highlighting edge differences. The incorporation of the rough edge map raised the IOU further but led to target sticking. In M5, SPE is replaced with DPE. We can see that prompt embedding at the tail of the model enhances target edge differences again and improves overall performance. Finally, BBM accurately eliminates false alarms and EGDSP achieves accurate annotation. In Figure 6, with the help of target energy initialization, our model quickly responds to the target area and accurately predicts the left rotor of the drone. In scene two, without double prompt embedding (Ours w/o DPE), neighboring targets blend. We further observe that the attention map also responded in the background region, revealing that the local shape prediction remains connected to the global context modeling. This finding underscores again the importance of preserving the entire scene.
Figure 6: Visual maps of ablation study on TEI and DPE. Incomplete labeling and adhesion are highlighted with arrows.
Downstream IRSTD
We used the inferred pseudo masks [^4] The number of data is 1505, training and test sets split as 8:2, we performed three cross-validations. from the test sets of SIRST, NUDT-SIRST, and IRSTD-1k to train the detection model, to validate the effectiveness of the pseudo labels for downstream IRSTD. Quantitative results are shown in Table 4. EDGSP outperforms other single-point prompt label generation methods. Although MCGC achieves the lowest ${{F}_{a}}$, it also has a low IoU score. Remarkably, EDGSP not only reduces time cost by 80% but also outperforms full labels in detection task. Even with coarse point prompts, It still achieves 99.5% of the performance of full labels. Visual results are shown in Figure 7, compared to full labels, DNA-Net trained with EDGSP’s pseudo mask demonstrates enhanced shape inference and lower false alarm. This reaffirms the applicability of the proposed methodology.
Figure 7: Visual results of DNA-Net. “Point Result” and “Full Result” indicate the outcomes of DNA-Net trained with the pseudo label from EDGSP and the full label, respectively. The correctly detected targets and false alarms are highlighted in orange and yellow circles, while incomplete shape inferences are marked with arrows.
Conclusion
In this article, we propose a novel annotation paradigm: enhancing detection networks with single-point prompts to accomplish the high-level IRSTLG. Specifically, we introduce the energy double guided single-point prompt (EDGSP) framework. It includes target energy initialization for effective shape evolution, double prompt embedding to prevent label adhesion, and bounding box-based matching to eliminate false alarms. This framework achieves accurate annotation on the SIRST, NUDT-SIRST, and IRSTD-1k datasets for the first time, and demonstrates superior performance compared to full labels in downstream IRSTD. In the future, we will focus on the coarse single-point prompt IRSTLG to further improve the utility of this framework.

### v4

@@
 Our intuition is that label generation requires just one more point prompt than target detection: IRSTLG can be regarded as an infrared small target detection (IRSTD) task with the target location hint.
-Based on this insight, we introduce an energy double guided single-point prompt (EDGSP) [^1] Code will be made public at https://github.com/xdFai.
+Based on this insight, we introduce an energy double guided single-point prompt (EDGSP) [^1] The code is available at https://github.com/xdFai/EDGSP.
 framework, which adeptly transforms the target detection network into a refined label generation method.
 Specifically, the proposed EDGSP includes: 1) target energy initialization (TEI) to create a foundational outline for sufficient shape evolution of pseudo label, 2) double prompt embedding (DPE) for rapid localization of interested regions and reinforcement of individual differences to avoid label adhesion, and 3) bounding box-based matching (BBM) to eliminate false alarms.
-Experimental results show that pseudo labels generated by three baselines equipped with EDGSP achieve 100% object-level probability of detection (${{P}_{d}}$) and 0% false-alarm rate (${{F}_{a}}$) on SIRST, NUDT-SIRST, and IRSTD-1k datasets, with a pixel-level intersection over union (IoU) improvement of 13.28% over state-of-the-art label generation methods.
-Additionally, the downstream detection task reveals that our centroid-annotated pseudo labels surpass full labels, even with coarse single-point annotations, it still achieves 99.5% performance of full labeling.
+Experimental results show that pseudo labels generated by three baselines equipped with EDGSP achieve 100% object-level probability of detection (${{P}_{d}}$) and 0% false-alarm rate (${{F}_{a}}$) on SIRST, NUDT-SIRST, and IRSTD-1k datasets, with a pixel-level intersection over union (IoU) improvement of 13.28% over state-of-the-art (SOTA) label generation methods.
+In the practical application of downstream IRSTD, EDGSP realizes, for the first time, a single-point generated pseudo mask beyond the full label.
+Even with coarse single-point annotations, it still achieves 99.5% performance of full labeling.
 Introduction
@@
 2024b).
-To isolate small targets from complex clutter, high-performance IRSTD methods typically extract discriminative features from extensive paired infrared images and pixel-level masks (Zhang et al.
+To isolate small targets from clutter, high-performance IRSTD methods typically extract discriminative features from extensive paired infrared images and pixel-level masks (Zhang et al.
 2022; Wu et al.
@@
 2024a).
-This learning paradigm highlights the need for large-scale, accurate mask annotation.
-However, infrared (IR) small targets typically occupy only a few pixels, exhibit a low signal-to-clutter ratio, and lack well-defined edges.
+This learning paradigm highlights the need for large-scale, accurate mask annotations.
+However, infrared (IR) small targets typically occupy only a few pixels and exhibit a low signal-to-clutter ratio and poorly defined edges.
 These characteristics render manual pixel-by-pixel annotation time-consuming and error-prone.
@@
 Single-point annotation [^2] Unless explicitly stated, centroid single-point is used.
-(SA) has recently demonstrated efficiency in labeling IR small targets, reducing the time needed by 80% compared to pixel-level annotation.
+(SA) has recently demonstrated efficiency in labeling IR small targets, reducing the time needed by 80% compared to pixel-level annotation (Li et al.
+2023a).
 As shown in Figure 1(a), LESPS (Ying et al.
@@
 Subsequently, a standardized label generation paradigm emerged.
-As illustrated in Figure 1(c), these pioneering works use SA to guide traditional model optimization for generating pseudo masks, such as MCLC (Li et al.
+As illustrated in Figure 1(c), these pioneering works use SA to guide traditional model optimization for generating pseudo labels, such as MCLC (Li et al.
 2023a), MCGC (Kou et al.
@@
 2024).
-Note that while these three works are claimed as “Single-Point Supervision”, they fundamentally belong to the label generation task with the single-point prompt as input.
+Note that while these three works are claimed as “single-point supervision”, they fundamentally belong to the label generation task with the “single-point prompt” as input.
 To avoid confusion, Figure 1(a)-(c) elucidates two critical links between these tasks.
@@
 2024).
-Experimental results indicate that existing IRSTLG methods fail to meet standards in AA and IoU.
-This issue arises because these traditional algorithms neglect to leverage target location cues to extract features from fully labeled data.
-Since well-designed IRSTD networks already achieve satisfactory target shape inference, and single-point prompts can be easily extracted from full labels, the primary goal of this paper is to design an effective “Upgrade Tool” that transforms the IRSTD network into a label generation method using the single-point prompt.
+Experimental results indicate that existing IRSTLG methods fail to meet standards in AA and IoU, primarily because these traditional algorithms neglect to leverage target location cues for feature extraction from fully labeled data.
+Since well-designed IRSTD networks already achieve satisfactory target shape inference, and single points can be easily extracted from full labels, the primary goal of this paper is to design an effective “upgrade tool” that transforms the IRSTD network into a more advanced label generation method under the single-point prompt.
 As shown in Figure 1(d), the essence of this paradigm is learning point prompt-based mask labeling experience from already labeled data.
@@
 A common approach involves using multiple extreme points or corners to precisely crop the target region, then pairing it with the encoded prompt and full labels for fully supervised training.
-Although this pipeline has yielded satisfactory results in routine areas, directly applying it to our task reveals three pitfalls.
+Although this pipeline has yielded satisfactory results in routine areas, directly applying it to our task exposes three pitfalls.
 1) Missing edge coordinates.
@@
 2023a).
-Two small targets in close proximity are easy to stick together and be mistakenly labeled as a single target, reducing ${{P}_{d}}$.
+Two small targets in close proximity can easily stick together and be mistakenly labeled as a single target, reducing ${{P}_{d}}$.
 3) Strong background clutter.
@@
 2024a).
-Even under effective location cues, this similarity still causes the deep model to misclassify pixels in these interference regions, resulting in false alarms.
-Figure 2: Comparing IRSTD and IRSTLG methods on mixed SIRST, NUDT-SIRST, and IRSTD-1k in $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{a}}({10}^{-6})$.
+Even with valid center-coordinate cues, this similarity still causes the deep model to misclassify pixels in interference regions, resulting in false alarms.
+Figure 2: Comparing IRSTD and IRSTLG methods on mixed datasets from SIRST, NUDT-SIRST, and IRSTD-1k in $IoU(\%)$, ${{P}_{d}}(\%)$, and ${{F}_{a}}({10}^{-6})$.
 EDGSP-Coa.
@@
 To tackle these problems, we propose a novel framework for IRSTLG, named energy double guided single-point prompt (EDGSP).
-Without requiring complex cropping, EDGSP enables end-to-end model target contours through the integration of local and global context provided by the point prompt and deep model.
-As shown in Figure 3, this framework consists of three key components.
+Without requiring complex cropping, EDGSP enables end-to-end modeling target contours through the integration of local and global context provided by the point prompt and deep network.
+As shown in Figure 3, this framework consists of three components.
 1) Target energy initialization (TEI): Inspired by IR targets that always have similar energy and shape distribution, we expand each point prompt into a Gaussian with local salience priors as the initial pattern of every target to evolve more complete outlines.
 2) Double prompt embedding (DPE): The encoded prompt is embedded twice.
-The first embedding ensures the model directly focuses on the interested region, even in shallow layers.
-The second embedding refines the target salience map, enhancing target boundaries and preventing shape merging.
-3) Bounding box-based matching (BBM): As a post-processing approach, we generate bounding boxes for each candidate target, then match the point prompt with each box by coordinates to eliminate clutter.
-Benefiting from the above structures, as shown in Figure 2, our annotation method showcases powerful shape prediction capabilities and achieves ${{P}_{d}}$=100%, ${{F}_{a}}$=0on three datasets.
+The first embedding ensures that the model directly focuses on the region of interest, even in shallow layers.
+The second embedding refines the target salience map, fine-tuning mask boundaries and preventing shape merging.
+3) Bounding box-based matching (BBM): As a post-processing step, we generate bounding boxes for each candidate target and then match the point prompt with each box by coordinates to eliminate clutter.
+Benefiting from the above structures, as shown in Figure 2, our method showcases powerful shape prediction capabilities and achieves ${{P}_{d}}$=100%, ${{F}_{a}}$=0on three datasets.
 The main contributions are summarised as follows.
   • To the best of our knowledge, we present the first study of the learning-based IRSTLG paradigm, and introduce EDGSP creating a crucial link between label generation and target detection task.
-  • We propose the target energy initialization (TEI), double prompt embedding (DPE), and bounding box-based matching (BBM) strategies to address the insufficient label shape evolution, label adhesion, and false alarm.
+  • We propose target energy initialization (TEI), double prompt embedding (DPE), and bounding box-based matching (BBM) strategies to address insufficient shape evolution, label adhesion, and false alarms.
   • For the first time, three baselines equipped with EDGSP achieve accurate annotation on three datasets.
@@
 To mitigate the semantic gap between encoders and decoders, DNA-Net (Li et al.
-2022) adopts dense nested interactive modules to facilitate progressive interaction of cross-scale features and adaptive enhancement of fine-grained features, while UIU-Net (Wu, Hong, and Chanussot 2022) directly nest multiple small U-Net into each codec of U-shaped structure.
+2022) adopts dense nested interactive modules to facilitate progressive interaction of cross-scale features and adaptive enhancement of fine-grained features, while UIU-Net (Wu, Hong, and Chanussot 2022) directly nests the small U-Net into each codec of the large U-shaped structure.
 To address the limitations in long-range information modeling, MTU-Net (Wu et al.
@@
 One working mode related to us is point-supervised segmentation.
-As a weak supervision form, point supervision only requires one or a few points as label (Qian et al.
+As a manner of weak supervision, point supervision only requires one or a few points as labels (Qian et al.
 2019; Zhao and Yin 2020; Li et al.
@@
 Close to our work, LESPS (Ying et al.
-2023) is the first to introduce single-point supervision to infrared small target detection task, based on mapping degradation phenomenon to guide point label evolution.
+2023) is the first to introduce single-point supervision to the infrared small target detection task, using the mapping degradation phenomenon to guide point label evolution.
 Unlike the aforementioned works, we still rely on fully annotated masks for supervision, yet employ single-point as a form of input to guide the IRSTD network.
-This approach reveals how to utilize additional prompts to achieve high-precision pseudo-label generation for small targets.
+This approach investigates how to utilize the additional prompt to achieve high-precision pseudo-label generation for IR small targets.
 Infrared Small Target Label Generation
 To the best of our knowledge, there are only three IRSTLG efforts.
-They are all based on the model-based single-point prompt methods.
+They are all model-based single-point prompt methods.
 As a pioneer, MCLC (Li et al.
@@
 2024).
-Recently, Li et al.
+Li et al.
 (Li et al.
-2024) designed a variational level set with an expected difference energy function to address the excessive regularization causing target contours to disappear.
-However, these works tend to treat the points as a ”Supervision” signal.
+2024) design a variational level set with an expected difference energy function to address the excessive regularization causing target contours to disappear.
+However, these works tend to treat the point as a ”Supervision” signal.
 This biased task orientation causes them to overlook the potential of combining point prompt-derived location cues to extract features from complete labeled datasets, culminating in pseudo labels with unclear edges, target adhesion, and false alarms.
 In this paper, we accurately define the operating paradigm where images and single-point prompts serve as inputs and produce pseudo labels as “single-point prompt for infrared small target label generation”.
-Unlike previous methods, we present the first study of a learning-based paradigm for IRSTLG, which artfully evolves the target detection network into an advanced label generation method with the aid of the single-point prompt.
+Unlike previous methods, we present the first study of a learning-based paradigm for IRSTLG, which artfully evolves the target detection network into a high-level label generation method with the aid of the single-point prompt.
 Figure 3: Illustration of the proposed energy double guided single-point prompt (EDGSP) framework.
@@
 Similarly, IOG (Zhang et al.
-2020) uses the approximate center of the target as the inside point and a set of symmetrical corner points as the outside points.
-After cropping the target of interest, the image, along with heat maps is fed into the segmentation model.
+2020) uses the approximate center of the target as the inside point and a set of symmetrical corner points as outside points.
+After cropping the target of interest, the image, along with the heatmap is fed into the segmentation model.
 Luo et al.
@@
 The proposed EDGSP differs from them in three ways: 1) It processes the entire image directly without cropping.
-2) It assigns explicit physical meaning to Gaussian heat maps, thereby using the double embedding strategy.
-3) EDGSP introduces a box-based matching approach specifically designed to eliminate false alarms.
+2) It assigns explicit physical meaning to Gaussian heatmap and introduces the double embedding strategy.
+3) EDGSP adopts a box-based matching approach customized to eliminate false alarms.
 Proposed Method
-The overview of our proposed EDGSP is shown in Figure 3. It is very concise and consists of three parts.
+The overview of our proposed EDGSP is shown in Figure 3. It is concise and consists of three parts.
 Target Energy Initialization
@@
 2013), we propose target energy initialization to predefine a basic profile for each target.
-As shown in Figure 3, given an input image ${\mathbf{{I}_{i}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$ and corresponding single-point prompt ${\mathbf{{I}_{p}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$ (Since the single-point prompt is a pixel within the target, it can be derived from the already labeled ground truth ${\mathbf{{G}_{t}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$).
+As shown in Figure 3, given an input IR image ${\mathbf{{I}_{i}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$ and corresponding single-point prompt ${\mathbf{{I}_{p}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$ (Since the single-point prompt is a pixel within the target, it can be derived from the already labeled ground truth ${\mathbf{{G}_{t}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$).
 2D Gaussian function is used to expand each point in ${\mathbf{{I}_{p}}}$ as an energy blob, thus creating a single init shape map ${\mathbf{{I}_{g}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$.
-This expansion simplifies the task by shifting from shaping area targets with points to refining contours using Gaussian blobs, aligning the prior distribution of the targets, and aiding the model in learning complete shape representations.
+This expansion simplifies the task by shifting from shaping area targets with points to refining contours using Gaussian blobs, aligning the prior distribution of the object, and aiding the model in learning complete shape representations.
 Double Prompt Embedding
@@
 where ${\mathcal{G}^{Conv}_{w/o}}$ stands for $\mathcal{G}$ that without the last convolution layer.
-In the second embedding, ${\mathbf{{I}_{g}}}$ is treated as a semantic map that reinforces the target quantity, it is concatenated with the ${\mathbf{L}}$ along the channel dimension to refine the model’s final mapping and prevent close-range target sticking.
+In the second embedding, ${\mathbf{{I}_{g}}}$ is treated as a semantic map that emphasizes the target quantity and edge difference.
+It is concatenated with the ${\mathbf{L}}$ along the channel dimension to refine the model’s final mapping and prevent close-range target sticking.
 ${\mathbf{{O}_{s}}}={Sigmoid}({Conv}(Concat({\mathbf{L}},{\mathbf{{I}_{g}}})))$ (3)
 where ${\mathbf{{O}_{s}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$ represents the saliency map, ${Conv}$ denotes the last layer of $\mathcal{G}$, typically constructed with a $1\times 1$ or $3\times 3$ convolution that squeezes channel and preserves the spatial dimensions.
-The binary cross entropy loss is then used to compute pixel-wise differences between the ${\mathbf{{O}_{s}}}$ and the ground truth ${\mathbf{{G}_{t}}}$.
-Table 1: Comparison with SOTA IRSTLG methods on SIRST, NUDT-SIRST, IRSTD-1k in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{at}}(\%)$, ${{F}_{a}}({10}^{-6})$.
+The binary cross entropy loss is then employed to compute pixel-wise differences between the ${\mathbf{{O}_{s}}}$ and the ground truth ${\mathbf{{G}_{t}}}$.
+Table 1: Comparison with SOTA IRSTLG methods on SIRST, NUDT-SIRST, and IRSTD-1k in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{at}}(\%)$, and ${{F}_{a}}({10}^{-6})$.
 The best and the second-best results are highlighted in bold and underlined, respectively.
@@
 Bounding Box-based Matching
-In previous works, eight connective regions matching is popular to eliminate false alarms during the label evolution process (Ying et al.
+In previous works, eight connective regions matching is commonly used to exclude false alarms during the label evolution process (Ying et al.
 2023; Kou et al.
 2024), ensuring that evolution information consistently applies to the initial point.
-However, this strategy ignores cases where the point prompt is located on or outside the target boundary, leading to incorrect processing results [^3] The shortcomings of this approach are detailed in the supplementary material..
+However, this strategy ignores cases where the point prompt is located on or outside the target boundary, leading to incorrect processing results [^3] The shortcomings of eight connective regions matching are detailed in the supplementary material..
 We tackle this issue by adding a compact design.
-In Figure 3(c), after performing eight-neighborhood clustering on the threshold-segmented candidate map ${\mathbf{{O}_{c}}}$, we generate bounding boxes for each component to create a b-box map ${\mathbf{{B}_{m}}}$.
+In Figure 3(c), after performing eight neighborhood clustering on the threshold-segmented candidate map ${\mathbf{{O}_{c}}}$, we generate bounding boxes for each component to create a b-box map ${\mathbf{{B}_{m}}}$.
 Then, match the point prompt ${\mathbf{{I}_{p}}}$ with ${\mathbf{{B}_{m}}}$.
@@
 Datasets.
-Experiments are conducted on three mixed datasets: SIRST (Dai et al.
+Following (Ying et al.
+2023), experiments are conducted on three mixed datasets: SIRST (Dai et al.
 2021), NUDT-SIRST (Li et al.
@@
 2022), which include 427, 1327, and 1001 infrared images with one or more targets, respectively.
-We used the method from (Li et al.
-2022) to partition the datasets for SIRST and NUDT-SIRST, and followed (Zhang et al.
-2022) for splitting IRSTD-1k.
+We use (Li et al.
+2022) to partition the datasets for SIRST and NUDT-SIRST, and follow (Zhang et al.
+2022) to split IRSTD-1k.
 Figure 4: Visual results of seven IRSTLG methods on the SIRST, NUDT-SIRST, IRSTD-1k.
@@
 2024) is used as the backbone of EDGSP.
-The peak half-width of the Gaussian is set to 4 and the segmentation threshold is set to 0.5.
+The peak half-width of the Gaussian is set to 4, and the segmentation threshold is set to 0.5.
 We compare all IRSTLG methods: MCLC (Li et al.
@@
 2023), using DNA-Net as the backbone, consistent with the original paper.
-We provide more details of experiments in the supplementary material.
-Table 2: The improvement brought by the EDGSP framework to seven SOTA IRSTD methods on SIRST, NUDT-SIRST, IRSTD-1k in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{at}}(\%)$, ${{F}_{a}}({10}^{-6})$, $Flops(G)$, and $Params(M)$.
+More details about the experiments are provided in the supplementary material.
+Table 2: The improvement brought by the EDGSP framework to seven SOTA IRSTD methods on mixed datasets from SIRST, NUDT-SIRST, and IRSTD-1k in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{at}}(\%)$, ${{F}_{a}}({10}^{-6})$, $Flops(G)$, and $Params(M)$.
 Method IRSTD Baseline IRSTD with EDGSP
@@
 As shown in Table 1, only EDGSP achieves ${{P}_{d}}$=100%, ${{F}_{a}}$=0%, ${{F}_{at}}$=0across three datasets at the same time.
-This indicates that our framework is currently the only one ensuring the most basic level of annotation quality.
+This indicates that our framework is currently the only one guaranteeing the most basic level of annotation quality.
 Additionally, EDGSP’s IoU surpasses the state-of-the-art (SOTA) labeling method (MCGC) by 14.27, demonstrating a superior shape prediction capability.
-Even with the coarse single-point prompt, EDGSP outperforms other methods with the centroid single-point prompt.
+Even with the coarse single-point prompt, EDGSP still outperforms other methods that use the centroid single-point prompt.
 Qualitative Results
-Figure 4 illustrates that the results of EDGSP are closer to the target silhouette in the infrared image than those of other models.
-In fact, many pseudo masks such as Misc_28, Misc_74, 000595, XDU104, and XDU697 even exceed the quality of ground truth (GT).
+Figure 4 illustrates that, compared to other models, the pseudo masks generated by EDGSP are closer to the target shape.
+In fact, many masks such as Misc_28, Misc_74, 000595, XDU104, and XDU697 even exceed the quality of ground truth (GT).
 We also observe that COM and MCLC tend to merge the target with the background, while LESPS, MCGC, and Robust SAM struggle to accurately infer the morphology of infrared small targets.
-Additionally, SAM is incapable of labeling infrared small targets.
-Figure 5: Quantitative results of three baselines at six cropping scales in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{at}}(\%)$.
+Additionally, SAM is unsuitable for annotating IR small targets.
+Figure 5: Quantitative results of three baselines at six cropping scales in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, and ${{F}_{at}}(\%)$.
 Dashed and solid lines represent label generation results from direct center-cropping and from using EDGSP on the cropped image, respectively.
@@
 Given that SCTransNet achieved the highest IoU score, we selected it as the backbone for EDGSP unless stated differently.
-Table 3: Based on SCTransNet (SCT), ablation study of the single prompt embedding (SPE), TEI, rough edge map (REM), DPE, and BBM on SIRST, NUDT-SIRST, and IRSTD-1k.
-SPE means directly concatenating the single-point prompt with the original image as input.
-More details of each module are discussed in the supplementary material.
+Table 3: Based on SCTransNet (SCT), ablation study of the single prompt embedding (SPE), TEI, rough edge map (REM), DPE, and BBM on mixed datasets from SIRST, NUDT-SIRST, and IRSTD-1k.
+SPE means concatenating the single-point prompt with the original image as input.
 Method SCT SPE TEI REM DPE BBM $IoU\uparrow$ ${{P}_{d}}\uparrow$ ${{F}_{at}}\downarrow$
@@
 With the crop size increasing, model performance initially improves and then declines, and fails to achieve high-performance labeling (${P}_{d}<100\%$, IoU$<$85).
-This is because small crop sizes fail to encompass large targets, while large crop sizes introduce excessive irrelevant padding at the image edges.
+This is because IR targets are multi-scale; small crop sizes fail to encompass large targets, while large crop sizes introduce excessive irrelevant padding at the image edges.
 Additionally, target overlap is challenging to manage when splicing the inferred image back into the original image.
 In summary, missing boundary coordinate information diminishes the effectiveness of the formal center-cropping strategy.
-Therefore, EDGSP directly processes the entire image.
-Table 4: Quantitative results of five IRSTD trained with five types of pseudo-labels and full labels on downstream test sets in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{a}}(\%)$.
+Therefore, our EDGSP directly processes the entire image.
+Table 4: Quantitative results of five IRSTD trained with five types of pseudo-labels and full labels on downstream test sets in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, and ${{F}_{a}}(\%)$.
 The best and the second-best results are highlighted in bold and underlined, respectively.
@@
 As shown in Table 3, SCTransNet is used as the baseline to validate the effectiveness of each proposed strategy.
-In M2, just incorporating single prompt embedding, the value of ${P}_{d}$ and IoU exceeds the full cropping strategies in Figure 5, illustrating the necessity of entire image processing.
-Using SPE, the IoU value further improved, and ${{P}_{d}}$ reached 100%, which demonstrated the SPE is capable of fleshing out target shapes and highlighting edge differences.
+In M2, just incorporating single prompt embedding, the value of ${P}_{d}$ and IoU exceeds the full cropping strategies in Figure 5, highlighting the need for whole image analysis.
+Using SPE, IoU further improved, and ${{P}_{d}}$ reached 100%, which demonstrated the SPE is capable of fleshing out target shapes.
 The incorporation of the rough edge map raised the IOU further but led to target sticking.
 In M5, SPE is replaced with DPE.
-We can see that prompt embedding at the tail of the model enhances target edge differences again and improves overall performance.
+We can see that prompt embedding at the tail of the model highlights target edge differences and improves overall performance.
 Finally, BBM accurately eliminates false alarms and EGDSP achieves accurate annotation.
-In Figure 6, with the help of target energy initialization, our model quickly responds to the target area and accurately predicts the left rotor of the drone.
-In scene two, without double prompt embedding (Ours w/o DPE), neighboring targets blend.
+Observe the visual maps in Figure 6.
+With the help of target energy initialization, our model quickly responds to the target area in the shallow layer and accurately predicts the left rotor of the drone.
+In scene two, without double prompt embedding (Ours w/o DPE), neighboring targets mix together.
 We further observe that the attention map also responded in the background region, revealing that the local shape prediction remains connected to the global context modeling.
 This finding underscores again the importance of preserving the entire scene.
+More details about each module are discussed in the supplementary material.
 Figure 6: Visual maps of ablation study on TEI and DPE. Incomplete labeling and adhesion are highlighted with arrows.
 Downstream IRSTD
-We used the inferred pseudo masks [^4] The number of data is 1505, training and test sets split as 8:2, we performed three cross-validations.
-from the test sets of SIRST, NUDT-SIRST, and IRSTD-1k to train the detection model, to validate the effectiveness of the pseudo labels for downstream IRSTD.
+We used the inferred pseudo masks from the test sets444 This dataset consists of 1079 samples, with the training and testing sets split 8:2.
+We performed three cross-validations.
+of SIRST, NUDT-SIRST, and IRSTD-1k to train the detection model, to validate the effectiveness of the pseudo labels for IRSTD.
 Quantitative results are shown in Table 4.
@@
 Remarkably, EDGSP not only reduces time cost by 80% but also outperforms full labels in detection task.
-Even with coarse point prompts, It still achieves 99.5% of the performance of full labels.
-Visual results are shown in Figure 7, compared to full labels, DNA-Net trained with EDGSP’s pseudo mask demonstrates enhanced shape inference and lower false alarm.
+Even with the coarse point prompt, EDGSP still achieves 99.5% performance of pixel-wise labeling.
+Visual results are shown in Figure 7.
+DNA-Net trained with EDGSP’s pseudo masks demonstrates enhanced shape and fewer false alarms compared to full labels.
 This reaffirms the applicability of the proposed methodology.
@@
 “Point Result” and “Full Result” indicate the outcomes of DNA-Net trained with the pseudo label from EDGSP and the full label, respectively.
-The correctly detected targets and false alarms are highlighted in orange and yellow circles, while incomplete shape inferences are marked with arrows.
+The correctly detected targets and false alarms are highlighted in orange and yellow circles, while incomplete shapes are marked with arrows.
 Conclusion
-In this article, we propose a novel annotation paradigm: enhancing detection networks with single-point prompts to accomplish the high-level IRSTLG.
-Specifically, we introduce the energy double guided single-point prompt (EDGSP) framework.
+In this paper, we propose a novel annotation paradigm: enhancing the IRSTD network with the single-point prompt to accomplish high-precision IRSTLG.
+Specifically, we introduce an EDGSP framework.
 It includes target energy initialization for effective shape evolution, double prompt embedding to prevent label adhesion, and bounding box-based matching to eliminate false alarms.
-This framework achieves accurate annotation on the SIRST, NUDT-SIRST, and IRSTD-1k datasets for the first time, and demonstrates superior performance compared to full labels in downstream IRSTD.
-In the future, we will focus on the coarse single-point prompt IRSTLG to further improve the utility of this framework.
+Three baselines equipped with EDGSP achieve accurate annotation on three IRSTD datasets for the first time.
+Excitingly, this is also the first time that a single point-generated pseudo mask outperforms full labels in practical IRSTD applications.
+In the future, it is worth investigating the coarse single-point prompt to further improve the utility of this framework.